C = 350 mil + 0.85Yd

I = 175 mil

G = 300 mil

X = 170 mil

Z = 120 mil

T = 0,15Y

Yf = R3100 million

Equilibrium Income: Y = C + I + G + (X - Z)

Disposable Income: Yd = (1 - t)Y


• What is the value of the marginal propensity to consume in this model?


• Calculate the Multiplier


1
Expert's answer
2022-03-24T16:12:33-0400

 "Y=C+I+G"


Y = C + I + G + (X - Z)

Y=350 mil + 0.85(1-0.15Y)+175 mil+ 300 mil+(170 mil-120 mil)

Y=350 mil + 0.85-0.1275Y +475 mil + 50 mil

Put like terms together

Y + 0.1275Y=0.85 + 525 mil

0.8725Y=0.85 + 525 mil